OpenStudy (anonymous):

How many terms are needed in the following geometric sequence to get within 10^-9 of zero? b.) i.) 4, -2, 1

OpenStudy (perl):

$$ \Large a_n = 4(-\frac{1}{2})^{n-1} $$

OpenStudy (perl):

so you need to solve $$ \Large {|4(- \frac{1}{2} ) ^{n-1} - 0 | \leq 10^{-9} } $$

OpenStudy (perl):

$$ \Large { |4(- \frac{1}{2} ) ^{n-1} - 0 | \leq 10^{-9} \\ \therefore \\ 4( \frac{1}{2} ) ^{n-1} \leq 10^{-9} \\ ~ \\ ( \frac{1}{2} ) ^{n-1} \leq \frac{10^{-9} }{4} \\~\\ \log ( \frac{1}{2} ) ^{n-1} \leq \log (\frac{10^{-9} }{4}) \\~\\ (n-1) \log (\frac{1}{2}) \leq \log (\frac{10^{-9} }{4}) \\~\\ (n-1) \leq \frac{\log (\frac{10^{-9} }{4})}{ \log (\frac{1}{2})} \\~\\ n \leq \frac{\log (\frac{10^{-9} }{4})}{ \log (\frac{1}{2})}+1 } $$